OpenStudy (anonymous):

Write an exponential function y=ab^2 for a graph that includes (2,24) and (3,48)

OpenStudy (anonymous):

plug and solve for a and b

OpenStudy (anonymous):

i think there is a mistake in the question \[ab^2=a\times b\times b\] is not an exponential function

OpenStudy (anonymous):

maybe it is \(y=a\times b^x\) ?

OpenStudy (anonymous):

ab^x sorry.

OpenStudy (anonymous):

ok then first you solve for \(b\) since you know \[a\times b^3=48\] and \[a\times b^2=24\] and \[\frac{a\times b^3}{a\times b^2}=b=\frac{48}{24}=2\] so you have \(b=2\)

OpenStudy (anonymous):

It was 24

OpenStudy (anonymous):

?

OpenStudy (anonymous):

now that we have \(b=2\) we know \(a\times 2^2=24\) so \(4a=24\) and therefore \(a=6\) and your function is \[y=6\times 2^x\]